You own a portfolio that is 25 percent invested in Stock X, 35 percent in Stock Y, and 40 percent in Stock Z. The expected returns on these three stocks are 10 percent, 13 percent, and 15 percent, respectively. What is the expected return on the portfolio?

PART A: is this a systematic risk or a unsystematic risk
Answer #1

Weight in Stock X, W1 = 0.25

Weight in Stock Y, W2 = 0.35

Weight in Stock Z, W3 = 0.40

Expected return on Stock X, R1 = 10%

Expected return on STock Y, R2 = 13%

Expected return on Stock Z, R3= 15%

Expected return on portfolio = W1 * R1 + W2 * R2 + W3 * R3 = 0.25* 0.1 + 0.35* 0.13 + 0.4* 0.15

= 0.13 = 13.05%

Systematic risk is market risk and cannot be diversified. In the above example, by creating a portfolio the unsystematic risk is reduced through diversification.